a wire of resistance 3 ohm is cut into three equal pieces ,which are joined to form a triangle .the equivalent resistance between any two corners of the triangle is ?


 

Grade:12

2 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Anju,

3 ohm wire cut into 3 equal pieces.........then each wire resistance is 1 ohm

then for the given arrangement: (2 resistors of 1 ohm each in series ) parallel with(1 ohm resistor)

2ohm in parallel with 1 ohm

2*1/(2+1)=2/3 ohm

charanendra bekalkar
14 years ago

since the resistance for a given material remains constant. THEREFORE out of three resistors, two are in series and the resultant is parellel with the remaining resistor

            3+3=6

            (6)(3)     = 2

             6+3

  therefore equivalent resistor is = 2 ohm